OpenStudy (anonymous):

How would you complete the square for 5x^2=2x+1

OpenStudy (fanduekisses):

ok so first of all move the 2x to the left side, so subtract 2x to both sides

OpenStudy (anonymous):

Okay, so then I would get 5x^2-2x=1?

OpenStudy (johnweldon1993):

Right^ now divide everything by 5 we want the 'x^2' term by itself

OpenStudy (anonymous):

And so If I do this then I will get x^2-2x/5=1/5 right?

OpenStudy (johnweldon1993):

Correct \[\large x^2 - \frac{2}{5}x = \frac{1}{5}\] So now, we take the coefficient of the 'x' term....and divide it by 2...we then square that result...what do we have from that?

OpenStudy (anonymous):

Well if I do this then I get x^2-1/5=1/5 but wouldn't I also multiply the -1/5 on the other side of the equation so then it would be x^2-1/5=1/25 because when completing the square whenever you add to one side of an equation you must also add to the other side then by squaring it I would get (x^2-1/5)^2=-1/25 and then I would square both sides and get x-1/5=the square root of negative 1 over 25? Or did I do something wrong?

OpenStudy (johnweldon1993):

Not quite! you didnt square the result originally \[\large x^2 - \frac{2}{5}x = \frac{1}{5}\] Divide 2/5 by 2 \[\large \frac{2}{5} \div 2 = \frac{1}{5}\] then square the result \[\large (\frac{1}{5})^2 = \frac{1}{25}\] and yes THIS you add to both sides of the equation \[\large x^2 - \frac{2}{5}x + \frac{1}{25} = \frac{1}{5} + \frac{1}{25}\] \[\large x^2 - \frac{2}{5}x + \frac{1}{25} = \frac{6}{25}\] Turn the left hand side into a square \[\large (x - \frac{1}{5})^2 = \frac{6}{25}\] Can you take it from there?

OpenStudy (anonymous):

Yes Thanks for the Help!!:)

OpenStudy (johnweldon1993):

Of course :)